# Why $w\neq-1$? Anthropic Selection in a $\Lambda$ + Axion Dark Energy Model

### v1

@@ Abstract
We study a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers can exist and observe an accelerating universe, we derive nontrivial constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below because extremely light axions would require the bare cosmological constant to be unnaturally close to zero to achieve accelerated expansion. As a result, the anthropically allowed axion mass range typically lies around $m=\mathcal{O}(10)\,H_{0}$ for a decay constant close to the Planck scale, where $H_{0}$ is the observed value of the Hubble constant. In this framework, the dark energy equation of state parameter $w_{0}$ generically deviates from −$-1$ by $\mathcal{O}(0.1)$, providing a natural explanation for why $w\neq-1$ may be expected. This outcome is intriguingly consistent with recent DESI hints of time-varying dark energy, and offers a compelling anthropic explanation within the $\Lambda$ + axion framework.
@@ I Introduction
Among the proposed solutions to the cosmological constant problem, the anthropic principle Weinberg (1987, 1989) has gained significant attention, especially in the context of the string landscape Susskind (2003); Polchinski (2006); Ellis and Smolin (2009). If the landscape contains a vast number of vacua with different vacuum energies, it is natural to expect that observers can only exist in regions of the multiverse where the cosmological constant falls within a narrow anthropic window: too large a value would suppress galaxy formation, while a negative value would lead to rapid recollapse. This line of reasoning offers a statistical explanation for the observed smallness of the cosmological constant.111 In fact, the observed amount of dark energy is smaller than the anthropic upper bound, which may call for explanation Sorini et al. (2024). This is beyond the scope of this letter.
@@
In this work, we go beyond the single-component picture in the anthropic argument and consider a scenario in which dark energy arises from the sum of a bare negative cosmological constant and the energy of an ultra-light axion. [^2] A related anthropic approach has been taken in the context of the strong CP problem, where the sum of a cosmological constant and an axion potential was considered as the effective cosmological constant Takahashi (2009); Kaloper (2019). Such axions are generically predicted in string theory, with masses spanning many orders of magnitude—a scenario often referred to as the axiverse Arvanitaki et al. (2010). If the axion is sufficiently light, it can remain displaced from its potential minimum over cosmological timescales and contribute to the present-day dark energy density.
@@
Our goal is not to determine the full anthropically allowed parameter space of the $\Lambda$ + axion dark energy model, which would be highly nontrivial due to the many free parameters and the complexity of possible cosmological histories. Instead, we restrict our attention to scenarios resembling our universe, where intelligent observers can exist at some point in cosmic history and observe cosmic acceleration. Within this context, we investigate what can be said about the properties of the axion and the dark energy equation-of-state parameter.
@@
In a universe where intelligent observers exist and observe an accelerating expansion, the total dark energy is required to lie below an anthropic upper bound on the cosmological constant. Although this condition does not directly constrain each component individually, we show that it nonetheless leads to nontrivial constraints on both the axion mass and the bare cosmological constant. Specifically, the axion mass is bounded from above because large masses would require the initial misalignment angle to be finely tuned near the hilltop in order to avoid early relaxation and recollapse. Conversely, the axion mass is bounded from below because, for too small masses, the bare cosmological constant must be tuned to be very close to zero to achieve accelerated expansion. These dual constraints result in a finite range of allowed axion masses and bare vacuum energies.
@@
Remarkably, within this constrained parameter space, the dark energy equation of state parameter $w$ is generically predicted to deviate from −$-1$ by $\mathcal{O}(0.1)$. This deviation arises naturally from the anthropic selection and is consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b), providing a compelling explanation within the $\Lambda$ + axion framework. See Refs. Tada and Terada (2024); Yin (2024); Cortês and Liddle (2024); Bhattacharya et al. (2024a); Mukherjee and Sen (2024); Notari et al. (2024); Jia et al. (2025); Hernández-Almada et al. (2024); Bhattacharya et al. (2024b); Berbig (2025); Wolf et al. (2025); Chakraborty et al. (2025); Borghetto et al. (2025); Nakagawa et al. (2025); Lee et al. (2025); Colgáin et al. (2025); Santos et al. (2025) for various explanations of the DESI results.
@@ II $\Lambda$ + Axion dark energy model
where $m$ is the axion mass, $f$ is its decay constant, and we choose the origin, $\phi=0$, to be the maximum of the potential. We set $f=M_{\mathrm{Pl}}$ throughout, but our results remain largely unchanged as long as $f$ is somewhat below the Planck mass (e.g., $f\sim 10^{17}\,{\rm GeV}$). We will return to the case with smaller decay constants in a later discussion. We also assume that the axion field is spatially homogeneous.
For our analysis, we treat $\rho_{\Lambda}$, $\theta_{\mathrm{i}}\equiv\phi_{\mathrm{i}}/f$, and $m$ as independent random variables with flat priors for $\rho_{\Lambda}$ and $\theta_{\mathrm{i}}$, while fixing all other parameters to the observed values. Here, $\phi_{\mathrm{i}}$ is the initial value of $\phi$.
In an expanding universe, the axion follows the equation of motion
@@
Anthropic considerations for a conventional cosmological constant impose an upper bound on its energy density, $\rho_{\Lambda}<\rho_{\mathrm{\Lambda}}^{\rm max}$, derived from the requirement that galaxies can form Weinberg (1987, 1989). In the following, we adopt the upper bound from the star formation history, $\rho_{\mathrm{\Lambda}}^{\rm max}=2306\rho_{\mathrm{DE,obs}}$ Sorini et al. (2024) with $\rho_{\mathrm{DE,obs}}=3M_{\mathrm{Pl}}^{2}H_{0}^{2}\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}\simeq 0.7$. A lower bound also follows from the requirement that the universe should not recollapse too soon after galaxy formation, allowing sufficient time for the development of intelligent life. In our scenario, the total dark energy density is given by the sum of a negative cosmological constant and the energy of an ultra-light axion, and it must lie within the following range:
$0<\rho_{\mathrm{DE,\mathrm{i}}}<\rho_{\mathrm{\Lambda}}^{\rm max},$ (6)
where $\rho_{\mathrm{DE,\mathrm{i}}}$ is the initial energy density of dark energy. We refer to this as a conditional anthropic bound. The lower bound reflects the need for the axion to uplift the negative vacuum energy and initially behave like a cosmological constant in order to drive accelerated expansion. The upper bound ensures that galaxy formation occurs before the axion starts oscillating. Furthermore, we impose the additional condition that the universe has not yet reached the turnaround point at the present time, to account for the time required after galaxy formation for intelligent life to emerge. This is a simplifying assumption, and our results remain essentially unchanged even if we adopt the less restrictive condition that the universe has not recollapsed by the present time or if we adopt a slightly different time $t_{0}$.
Since we fix $f=M_{\mathrm{Pl}}$, the Hubble parameter is always smaller than the axion mass $m$ during the dark-energy-dominated epoch. As a result, if $m$ is much larger than the Hubble parameter, the axion would have rolled down the potential by now, and the universe would have started to contract and recollapse unless the initial field value were sufficiently close to the hilltop. This degree of fine-tuning becomes increasingly severe for larger values of $|\rho_{\Lambda}|$, as a larger axion mass is required to uplift the total vacuum energy into the conditional anthropic window (6).
@@
The equation-of-state parameter at the present time $t=t_{0}$ is given by
@@
We evaluate the evolution of the axion and $w_{0}$ both numerically and analytically. In Fig. 1, we show the parameter region allowed by the conditional anthropic bound for $m=100H_{0}$, $10H_{0}$, and $H_{0}$. We fix the range of $\rho_{\Lambda}$ following the condition (6). Note that the scales of the horizontal and vertical axes are different in the three panels. In the gray-shaded regions, the universe starts to contract before the present time, and we exclude these regions based on the requirement that the universe must not recollapse too soon after galaxy formation. We also show $w_{0}$ by the color shading for the allowed regions. Smaller values of $\theta_{\mathrm{i}}$ lead to smaller $w_{0}$, as the axion remains closer to the hilltop. The required tuning of $\theta_{\mathrm{i}}$ becomes more severe for small $\rho_{\Lambda}$ because the Hubble parameter is small, which allows the axion to roll down the potential more easily. Near the boundaries of the allowed region, $w_{0}$ becomes larger since the denominator of Eq. (17) becomes smaller. In the narrow regions at the edge of the allowed regions, the sign of $w_{0}$ flips due to the negative value of the denominator. While this is visible for $m=H_{0}$, such a region cannot be resolved for $m=100H_{0}$ due to the limited resolution of the plot.
@@
Figure 1: Colored regions indicate the predicted values of $w_{0}$, and dashed lines show analytic estimates of the boundaries between them. The axion mass is taken to be $m=100H_{0}$, $10H_{0}$, and $H_{0}$ from top to bottom. The gray-shaded region in each panel is excluded because the universe recollapses before the present time.
@@
We also show the results of the analytic estimate with dashed lines. Below the black dashed lines, $\rho_{m,0}+\rho_{\Lambda}+V(\phi(t_{0}))<0$ or $\phi(t_{0})>\pi f$ holds. Note that we excluded the kinetic energy from the total energy in the criterion for recollapse because it can be wrongly large for $\phi(t_{0})\gtrsim f$ in the analytic approximation. The colored dashed lines correspond to $w_{0}=-0.9,-0.6,-0.3,0$ basically from left to right. For $m=H_{0}$, there also appears an isolated region with $w<-0.9$ for the reason mentioned above. While the analytic estimate of $w_{0}$ agrees well with the numerical results for small $w_{0}$ and large $m$, it worsens in the opposite case, where the approximations of $\phi\ll f$ and $\rho_{\phi}=\mathrm{const}.$ break down.
@@
With the flat priors for $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$, we can evaluate the likelihood of $\ln m$ by integrating the area of the colored regions in Fig. 1. We show the likelihood of $\ln m$ in Fig. 2. The vertical axis is normalized so that the maximum value equals unity. We find a peak at $m=m_{\mathrm{peak}}\simeq 49.2H_{0}$, which is determined by $2m^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. This behavior can be understood as follows. For small $m$, the allowed value for $\rho_{\Lambda}$ is limited by −$-2m^{2}f^{2}<\rho_{\Lambda}<0$ by Eq. (6), and the likelihood increases roughly ∝$\propto m^{2}$. For large $m$, while the interval of the allowed $\rho_{\Lambda}$ becomes equal to $\rho_{\mathrm{\Lambda}}^{\mathrm{max}}$, independent of $m$, the required tuning of $\theta_{\mathrm{i}}$ near the hilltop becomes exponentially severe, suppressing the likelihood. While the peak position of the axion mass distribution depends on the values of $\rho_{\Lambda}^{\mathrm{max}}$ and $f$ as $2m_{\mathrm{peak}}^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$, the existence of the sharp peak itself does not depend on their values.
@@
In Fig. 3, we show the probability distribution of $w_{0}$ for fixed $m$. Here, we sample $w_{0}$ from −$-1$ to $1$ in increments of $0.1$, and $w_{0}>1$ are collected into the rightmost bin. We evaluate the probability by integrating the area with corresponding values of $w_{0}$ in Fig. 1 and normalize it such that the total equals unity. We find that while the most likely value of $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant. Thus, our results suggest that, in a universe where intelligent observers can observe the accelerated expansion, the dark energy equation-of-state parameter is likely to deviate from −$-1$ at the level of ${\cal O}(0.1)$.

### v2

@@ Abstract
We study a dark energy model composed of a bare negative cosmological constant and a single ultra-light axion, motivated by the string axiverse. Assuming that intelligent observers arise and observe, as in our universe, the onset of dark-energy-driven acceleration following matter domination, and that this acceleration persists to the present, we derive nontrivial constraints on both the axion mass and the bare cosmological constant. The axion mass is bounded from above to avoid fine-tuning of the initial misalignment angle near the hilltop, and from below because too light axions cannot achieve accelerated expansion due to their limited energy budget. As a result, the anthropically allowed axion mass range typically lies around $m=\mathcal{O}(10)\,H_{0}$ for a decay constant close to the Planck scale, where $H_{0}$ is the observed value of the Hubble constant. In this framework, the dark energy equation-of-state parameter $w_{0}$ generically deviates from −$-1$ by $\mathcal{O}(0.1)$, providing a natural explanation for why $w\neq-1$ may be expected. We also find that, for a decay constant slightly smaller than the Planck scale, the peak value of dark energy density is significantly smaller than the anthropic bound on the cosmological constant and can be close to the observed value. These outcomes are intriguingly consistent with recent DESI hints of time-varying dark energy, and offer a compelling anthropic explanation within the $\Lambda$ + axion framework.
††preprint: TU-1263
@@ I Introduction
Among the proposed solutions to the cosmological constant problem, the anthropic principle Weinberg (1987, 1989) has gained significant attention, especially in the context of the string landscape Susskind (2003); Polchinski (2006); Ellis and Smolin (2009). If the landscape contains a vast number of vacua with different vacuum energies, it is natural to expect that observers can only exist in regions of the multiverse where the cosmological constant falls within a narrow anthropic window: too large a value would suppress galaxy formation, while a negative value would lead to rapid recollapse. This line of reasoning offers a statistical explanation for the observed smallness of the cosmological constant.111 In fact, the observed amount of dark energy is smaller than the anthropic upper bound by about three orders of magnitude, which may call for an explanation Sorini et al. (2024). See Appendix A for a possible resolution within our framework.
@@
In this work, we go beyond the single-component picture in the anthropic argument and consider a scenario in which dark energy arises from the sum of a bare negative cosmological constant and the energy of an ultra-light axion. [^2] A related anthropic approach has been taken in the context of the strong CP problem, where the sum of a cosmological constant and an axion potential was considered as the effective cosmological constant Takahashi (2009); Kaloper (2019). Such ultra-light axions are generically predicted in string theory, where a large number of axions with masses spanning many orders of magnitude are expected—a scenario often referred to as the axiverse Arvanitaki et al. (2010). If the axion is sufficiently light, it can remain displaced from its potential minimum over cosmological timescales and contribute to the present-day dark energy density Svrcek (2006). The explanation of dark energy with a negative cosmological constant and an ultra-light axion has been explored in Refs. Ruchika et al. (2023); Luu et al. (2025). See also Refs. Cardenas et al. (2003); Dutta et al. (2020); Visinelli et al. (2019); Calderón et al. (2021); Sen et al. (2022); Adil et al. (2023); Menci et al. (2024) for related works.
@@
Our goal is not to determine the full anthropically allowed parameter space of the $\Lambda$ + axion dark energy model, which would be highly nontrivial due to the many free parameters and the complexity of possible cosmological histories. Instead, we restrict our attention to universes resembling our own, where intelligent observers arise and observe that the universe began accelerating due to dark energy after matter domination, and that this acceleration still continues today. Within this context, we investigate what can be said about the properties of the axion, the dark energy equation-of-state parameter, and the dark energy density.
@@
Under this restricted anthropic condition, the total dark energy is required to lie below an anthropic upper bound on the cosmological constant. Although this condition does not directly constrain each component individually, we show that it nonetheless leads to nontrivial constraints on both the axion mass and the bare cosmological constant. Specifically, the axion mass is bounded from above because large masses would require the initial misalignment angle to be finely tuned near the hilltop in order to avoid early relaxation and recollapse. Conversely, the axion mass is bounded from below because, for too small masses, the axion cannot achieve accelerated expansion due to its small energy budget. These dual constraints result in a finite range of allowed axion masses and bare vacuum energies.
@@
Remarkably, within this constrained parameter space, the dark energy equation-of-state parameter $w$ is generically predicted to deviate from −$-1$ by $\mathcal{O}(0.1)$. This deviation arises naturally from the anthropic selection and is consistent with recent DESI results hinting at time-varying dark energy Adame et al. (2025); Abdul Karim et al. (2025a, b), providing a compelling explanation within the $\Lambda$ + axion framework. See Refs. Tada and Terada (2024); Yin (2024); Cortês and Liddle (2024); Bhattacharya et al. (2024a); Mukherjee and Sen (2024); Notari et al. (2024); Jia et al. (2025); Hernández-Almada et al. (2024); Bhattacharya et al. (2024b); Berbig (2025); Wolf et al. (2025); Chakraborty et al. (2025); Borghetto et al. (2025); Nakagawa et al. (2025); Lee et al. (2025); Colgáin et al. (2025); Santos et al. (2025) for various explanations of the DESI results.
@@ II $\Lambda$ + Axion dark energy model
where $m$ is the axion mass, $f$ is its decay constant, and we choose the origin, $\phi=0$, to be the maximum of the potential. We set $f=M_{\mathrm{Pl}}$ throughout the main text, and we will return to the case with smaller decay constants in a later discussion. We also assume that the axion field is spatially homogeneous.
For our analysis, we treat $\rho_{\Lambda}$, $\theta_{\mathrm{i}}\equiv\phi_{\mathrm{i}}/f$, and $m$ as independent random variables with flat priors for $\rho_{\Lambda}$, $\theta_{\mathrm{i}}$, and $\ln m$, while fixing all other parameters to the observed values. Here, $\phi_{\mathrm{i}}$ is the initial value of $\phi$.
In an expanding universe, the axion follows the equation of motion,
@@
Anthropic considerations for a conventional cosmological constant alone impose an upper bound on its energy density, $\rho_{\Lambda}<\rho_{\mathrm{\Lambda}}^{\rm max}$, derived from the requirement that galaxies can form Weinberg (1987, 1989). In the following, we adopt the upper bound from the star formation history, $\rho_{\mathrm{\Lambda}}^{\rm max}=2306\rho_{\mathrm{DE,obs}}$ Sorini et al. (2024) with $\rho_{\mathrm{DE,obs}}=3M_{\mathrm{Pl}}^{2}H_{0}^{2}\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}\simeq 0.7$. A lower bound also follows from the requirement that the universe should not recollapse too soon after galaxy formation, allowing sufficient time for the development of intelligent life.
In this work on the axion dark energy, we further require that intelligent life observe accelerated expansion of the universe that persists to the present epoch. The accelerated expansion is realized when
$w_{\mathrm{tot}}\equiv\frac{p_{\mathrm{DE}}}{\rho_{m}+\rho_{\mathrm{DE}}}<-\frac{1}{3}\ ,$ (6)
where $p_{\mathrm{DE}}=\dot{\phi}^{2}/2-V(\phi)-\rho_{\Lambda}$ is the pressure of the dark energy component. In our scenario, the total dark energy component is given by the sum of a negative cosmological constant and an ultra-light axion and it satisfies $p_{\mathrm{DE}}\geq-\rho_{\mathrm{DE}}$. Thus, $\rho_{\mathrm{DE}}$ must become larger than $\rho_{m}/2$ at some point before the current time to realize $w_{\mathrm{tot}}<-1/3$. As a result, we obtain the condition for $\rho_{\mathrm{DE,i}}$ to satisfy our requirement as
$\frac{\rho_{m,0}}{2}<\rho_{\mathrm{DE,\mathrm{i}}}\lesssim\rho_{\mathrm{\Lambda}}^{\rm max},$ (7)
where $\rho_{\mathrm{DE,\mathrm{i}}}$ is the initial energy density of dark energy. We refer to this as a conditional anthropic bound. The lower bound reflects the need for the axion to uplift the negative vacuum energy and initially behave like a cosmological constant in order to drive accelerated expansion. The upper bound ensures that galaxy formation occurs before the axion starts oscillating. Note that the upper bound, $\rho_{\Lambda}^{\mathrm{max}}$, is derived under the assumption that the dark energy component solely consists of a cosmological constant. In our scenario, however, the axion should remain around the hilltop for accelerated expansion to persist until now, and so the dark energy density is expected to evolve only slowly. Therefore, we adopt $\rho_{\Lambda}^{\mathrm{max}}$ as the effective upper bound on the initial value of the total dark energy density.
Since we fix $f=M_{\mathrm{Pl}}$, the Hubble parameter is always smaller than the axion mass $m$ during the dark-energy-dominated epoch. In particular, if $m$ is much larger than the Hubble parameter, the axion would have rolled down the potential by now, and the universe would have ceased accelerating unless the initial field value were sufficiently close to the hilltop. This degree of fine-tuning becomes increasingly severe for larger values of $|\rho_{\Lambda}|$, as a larger axion mass is required to uplift the total vacuum energy into the conditional anthropic window (7).
@@
The dark energy equation-of-state parameter at the present time $t=t_{0}$ is given by
@@
We evaluate the evolution of the axion and $w_{0}$ both numerically and analytically. In Fig. 1, we show the parameter region allowed by the conditional anthropic bound for $m=100H_{0}$, $10H_{0}$, and $H_{0}$. We fix the range of $\rho_{\Lambda}$ following the condition (7). Note that the scales of the horizontal and vertical axes are different in the three panels. In the gray-shaded regions, our requirements are not satisfied: the universe never experiences accelerated expansion, or the accelerated expansion does not last until the current time. We also show $w_{0}$ by the color shading for the allowed regions. Smaller values of $\theta_{\mathrm{i}}$ lead to smaller $w_{0}$, as the axion remains closer to the hilltop. The required tuning of $\theta_{\mathrm{i}}$ becomes more severe for small $\rho_{\Lambda}$ (i.e., large $|\rho_{\Lambda}|$) because the Hubble parameter is small, which allows the axion to roll down the potential more easily. For $m=H_{0}$, $w_{0}$ cannot deviate significantly from −$-1$, namely, the axion should remain around the hilltop at the current time, since otherwise the accelerated universe cannot be achieved due to the small energy budget of the axion.
@@
Figure 1: Colored regions indicate the predicted values of $w_{0}$, and dashed lines show analytic estimates of the boundaries between them. The axion mass is taken to be $m=100H_{0}$, $10H_{0}$, and $H_{0}$ from top to bottom. The gray-shaded region in each panel is excluded due to the requirement for the accelerated expansion to persist until the current time.
@@
We also show the results of the analytic estimate with dashed lines. Below the black dashed lines, $\rho_{\mathrm{DE,i}}<\rho_{m,0}/2$ or $w_{0}>-1/3$ holds. The colored dashed lines correspond to $w_{0}=-0.9,-0.7,-0.5$ from left to right. For $m=H_{0}$, the black dashed line overlaps with the colored lines. While the analytic estimate of $w_{0}$ agrees well with the numerical results for small $w_{0}$ and large $m$, it worsens in the opposite case, where the approximations of $\phi\ll f$ and $\rho_{\phi}=\mathrm{const}.$ break down.
@@
With the flat priors for $\theta_{\mathrm{i}}$ and $\rho_{\Lambda}$, we can evaluate the likelihood of $\ln m$ by integrating the area of the colored regions in Fig. 1 over a denser sampling of the mass. We show the likelihood of $\ln m$ in Fig. 2. The vertical axis is normalized so that the maximum value equals unity. We find a peak at $m=m_{\mathrm{peak}}\simeq 49.2H_{0}$, which is determined by $2m^{2}f^{2}=\rho_{\Lambda}^{\mathrm{max}}$. This behavior can be understood as follows. Assuming that the axion stays near the hilltop, the allowed range of $\rho_{\Lambda}$ is limited as −$-2m^{2}f^{2}+\rho_{m,0}/2\lesssim\rho_{\Lambda}\lesssim{\rm Min}[0,\rho_{\Lambda}^{\rm max}-2m^{2}f^{2}]$. For small $m$ satisfying $\rho_{m,0}/2\ll 2m^{2}f^{2}<\rho_{\Lambda}^{\rm max}$, the interval of $\rho_{\Lambda}$ is approximately proportional to $m^{2}$. So, the likelihood increases roughly ∝$\propto m^{2}$. For large $m$ satisfying $2m^{2}f^{2}>\rho_{\Lambda}^{\rm max}$, while the interval of the allowed $\rho_{\Lambda}$ becomes approximately equal to $\rho_{\mathrm{\Lambda}}^{\mathrm{max}}$, independent of $m$, the required tuning of $\theta_{\mathrm{i}}$ near the hilltop becomes exponentially severe, suppressing the likelihood. Note that the peak position of the axion mass distribution depends on $\rho_{\Lambda}^{\mathrm{max}}$ as $m_{\mathrm{peak}}^{2}=\rho_{\Lambda}^{\mathrm{max}}/(2f^{2})$. In particular, if there exists some unknown anthropic condition that imposes a much tighter upper bound on $\rho_{\Lambda}$, which is close to the observed dark energy density, the peak of likelihood comes to $m_{\mathrm{peak}}=\mathcal{O}(H_{0})$. Moreover, this estimate for $m_{\mathrm{peak}}$ is further corrected for smaller $f$. See Appendix A for the different choices of $\rho_{\Lambda}^{\mathrm{max}}$ and $f$.
@@
In Fig. 3, we show the probability distribution of $w_{0}$ for fixed $m$. Here, we sample $w_{0}$ from −$-1$ to −$-0.3$ in increments of $0.05$. Note that the accelerated expansion at the current time implies $w_{0}<-1/3$. We evaluate the probability by integrating the area with corresponding values of $w_{0}$ in Fig. 1 and normalize it such that the total equals unity. We find that while the most likely value of $w_{0}$ lies close to −$-1$, the deviation by ${\cal O}(0.1)$ is statistically significant especially for $m=100H_{0}$ and $10H_{0}$. Given that the mass likelihood of $\ln m$ has a peak at $10H_{0}<m<100H_{0}$, our results suggest that, in a universe where intelligent observers can observe the accelerated expansion, the dark energy equation-of-state parameter is likely to deviate from −$-1$ at the level of ${\cal O}(0.1)$.
@@
Finally, we show the probability distribution of $\rho_{\mathrm{DE,i}}$ with the flat priors for $\theta_{\mathrm{i}}$, $\rho_{\Lambda}$, and $\ln m$ in Fig. 4. In a similar way to Fig. 3, we integrate the area of the colored regions with $\rho_{\mathrm{DE,i}}$ in each bin for different values of $\ln m$ and sum the area over $\ln m$ sampled with a uniform interval. Here, we take the bin for $24000<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<$ 2400 in increment of $100$. Note that the rightmost bin corresponds to $23062300<\rho_{\mathrm{DE,i}}/\rho_{\mathrm{DE,obs}}<$ 2306 due to the requirement (7) and thus has a lower probability. The probability distribution for $\rho_{\mathrm{DE,i}}$ spreads out over the region allowed by Eq. (7), and thus we cannot account for the hierarchy between $\rho_{\mathrm{DE,obs}}$ and $\rho_{\Lambda}^{\mathrm{max}}$ in this case. On the other hand, as we discuss in Appendix A, $\rho_{\mathrm{DE,i}}$ close to the observed value is favored when we choose a smaller $f$ such as $f=0.1M_{\mathrm{Pl}}$. Thus, for $f={\cal O}(0.1)M_{\mathrm{Pl}}$, $w_{0}$ naturally deviates from −$-1$ by ${\cal O}(0.1)$, and the dark energy density approaches the observed value, resolving the hierarchy between $\rho_{\rm DE,0}$ and $\rho_{\Lambda}^{\rm max}$.
Figure 4: Histogram of $\rho_{\mathrm{DE,i}}$ marginalized over $\theta_{\mathrm{i}}$, $\rho_{\Lambda}$, and $\ln m$.